Wildhorse Corporation operates a retail computer store. To improve delivery services to customers, the company purchases four new trucks on April 1, 2017. The terms of acquisition for each truck are described below.

1.
Truck #1 has a list price of $15,150 and is acquired for a cash payment of $14,039.
2.
Truck #2 has a list price of $16,160 and is acquired for a down payment of $2,020 cash and a zero-interest-bearing note with a face amount of $14,140. The note is due April 1, 2018. Wildhorse would normally have to pay interest at a rate of 9% for such a borrowing, and the dealership has an incremental borrowing rate of 8%.
3.
Truck #3 has a list price of $16,160. It is acquired in exchange for a computer system that Wildhorse carries in inventory. The computer system cost $12,120 and is normally sold by Wildhorse for $15,352. Wildhorse uses a perpetual inventory system.
4.
Truck #4 has a list price of $14,140. It is acquired in exchange for 930 shares of common stock in Wildhorse Corporation. The stock has a par value per share of $10 and a market price of $13 per share.


Prepare the appropriate journal entries for the above transactions for Wildhorse Corporation. (Round present value factors to 5 decimal places, e.g. 0.52587 and final answers to 0 decimal places, e.g. 5,275. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

No.

Account Titles and Explanation

Debit

Credit

1.


2.




3.




4.
Answer #1
Sno. Accounts and explanation Debit Credit
1. Truck # 1 $14,039
Cash $14039
(Purchased truck 1 for cash)
2. Truck # 2 ($14,140 x 0.91743) + $2020 [Note 1] $14,992
Discount on notes payable [ bal fig.] $1,168
Cash $2,020
Notes payable $14,140
(Purchased truck 2 for cash and notes payable)
3. Truck # 3 $15,352
Cost of goods sold $12,120
Inventory $12,120
Sales Revenue $15,352
(Purchased truck 3 in exchange of computer system)
4. Truck # 4 (930shares x $13) $12,090
Common stock (930 x $10) $9,300
Paid in capital in excess of par (930 x $3) $2,790
Purchased Truck 4 in exchange for 930shares of common stock)

Note 1:

Truck # 2 is worth the down payment + the Present Value of the note. Using the market rate of 9%, we find the PV of the note. Add that to the down payment to get the truck value.

PV factor (9%,1year) = 1 / (1+0.09) = 1 / 1.09 = 0.91743

Truck value = ($14,140 x 0.91743) + $2020 = $14,992